a) With the aid of a sketch explain the method of determining the dryness fraction of steam using a separating and throttling calorimeter,

(9 Marks)

b) The following observations were taken with a separating and a throttling

calorimeter arranged in series

Water separated = 2 kg. steam discharged from the throttling calorimeter = 20.5

kg temperature of steam after thretting 110
1
Expert's answer
2021-08-03T04:07:13-0400

A combined separating and throttling calorimeter is used to remove the drawback of Separating Calorimeter (i.e. water particles from wet steam are not fully separated) and the drawback of Throttling Calorimeter (i.e. not suitable for very wet steam).


Principle: In this calorimeter, the moisture from wet steam sample is first removed in separating calorimeter, so that dryness fraction of wet steam sample is increased above 0.95 before steam sample is entering into throttling calorimeter. During this process pressure and temperature remains constant.

The sample coming out from separating calorimeter is then passed through throttling calorimeter where it expands to superheated steam.


Material: It consists of a combined separating calorimeter and throttling calorimeter as discussed in previous sections.


Procedure: A sample of wet steam from main steam pipe (represented by point ‘A’) is first enters the separating calorimeter through a sampling tube. While passing through separating calorimeter, most of the moisture from wet steam is separated out and collected in the inner chamber of the separating calorimeter where it is measure with scale. Now, comparatively dry steam having dryness fraction above 0.95 from separating calorimeter (represented by point ‘B’) is passed into throttling calorimeter through throttling valve where it expands to superheated steam (represented by point ‘C’). The superheated steam after passing through throttling calorimeter is then exhausted from throttling calorimeter into a condenser, provided at the bottom of the throttling calorimeter, where it is finally condensed and measured.


The representation of separating-throttling processes of combined separating and throttling calorimeter on T-s and h-s diagrams is shown in Fig.

During steam sampling: Note down the following observations:

Pressure before throttling = p1 (say bar)

Pressure after throttling = p2 (say bar)

Temperature of a throttled superheated steam = t (say °C)


After steam sampling: Note down the following observations:

Mass of condensate collected in the container provided at the bottom of throttling calorimeter (i.e. mass of steam coming from separating calorimeter into throttling calorimeter through throttling valve) = M (say kg)


Mass of water collected in the inner chamber of the separating calorimeter (i.e. mass of water separated in the separating calorimeter = m (say kg)





Combined Separating and Throttling Calorimeter


Calculations of dryness fraction of wet steam, x:


Partial dryness fraction of wet steam coming from main steam pipe, measured by separating calorimeter alone.


Dryness fraction of steam entering into throttling calorimeter from separating calorimeter through throttling valve, measured by throttling calorimeter alone (by using equation 26.3)


Where,

hf,B is enthalpy of saturated water at p1 (from steam tables) , kJ/kg.

hfg,B is Latent heat of vapouization at p1 (from steam tables),kJ/kg

hsup,C is enthalpy of a throttled superheated steam corresponding to pressure p2 and temperature tsup,C (from superheated steam table),kJ/kg .


Mass of dry steam in the wet steam sample collected from main steam pipe = the mass of dry steam entering the throttling calorimeter = x2 . M


Finally, The total dryness fraction of wet steam collected from main steam pipe measure by Combined Separating and throttling Calorimeter

Dryness fraction = x1.x2


This shows that the dryness fraction of steam measured from combined separating and throttling calorimeter is the product of dryness fraction x1 and x2 measured by separating and throttling calorimeter respectively.